how do i find the solution for 5/4x+1/8x=9/8+x solve for x
Answer by stanbon(75887)

how do i find the solution for
(5/4)x+(1/8)x = (9/8)+x solve for x
----------
Multiply thru by 8 to get rid of the fractions:
5x + x = 9 + 8x
----
6x = 9 + 8x
2x = -9
x = -9/2
==============
Cheers,
Stan H.